OpenStudy (anonymous):

Find the inverse of each function: f(x)=x+2/3

OpenStudy (anonymous):

f(x)= -1/x +-3/2 (i think)

OpenStudy (anonymous):

It's supposed to be f-1(x) = 3x-2

OpenStudy (anonymous):

oopsie doodle came to town :P

OpenStudy (anonymous):

y = x + 2/3 Solve for x: x = y - 2/3 Interchange x and y f^-1 (x) = x - 2/3